A cylinder of mass M and radius R is resting on a horizontal platform(which is parallel to the x-y plane)with its axis fixed along the y axis and free to rotate about its axis. The platform is given a motion in the x direction given by x=A cos(ωt).there is no slipping between the cylinder and the platform.The max torque acting on the cylinder during its motion is?

Grade:11

2 Answers

Chetan Mandayam Nayakar
12 years ago

Dear Debadutta,

acceleration a=d2x/dt2=-w2Acos(wt),lamaxl=w2A,

maximum angular acceleration=w2A/R

Torque=Iα=(MR2/2)(w2A/R)=(MRAw2/2)

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Debadutta !!!

 


Regards,

Askiitians Experts

CHETAN MANDAYAM NAYAKAR

Ashwin Muralidharan IIT Madras
12 years ago

Hi Debadutta,

 

The max acn of the platform would be:

2.

 

So max Torque, ζmax = A(MRω2)/2.

 

Best Regards,

Ashwin (IIT Madras).